OpenStudy (anonymous):

If x varies inversely as y and directly as t, and x = 12 when t = 10 and y = 25, find y when x is 6 and t = 3

OpenStudy (anonymous):

x =kt/y 12 = 10k/25 275 =10k k = 27.5 6 = 3(27.5)/y y = 3(27.5)/6 y = 13.75

OpenStudy (anonymous):